# A Cartesian Cut-Cell Two-Fluid Method for Two-Phase Diffusion Problems

### v1

@@ 4.1.2 Gradient operator
The interfacial contribution (last term in (32)) is approximated by
@@
Finally, the constitutive relation (4) is imposed in discrete form by
@@ 4.2 Semi-discrete interface conditions
The interfacial conservation law (7) corresponds to continuity of the normal diffusive flux,
@@
The integrals in Eq. (35) correspond to the interfacial contributions appearing in the Gauss decomposition in Eq. (26). They are approximated using the discrete interfacial divergence operators $\operatorname{div}^{\alpha\gamma}$ introduced in Eq. (30). Consequently, the flux balance is discretized as
@@
For mixed cells $(i,j)$, the jump/closure condition in Eq. (8) (or in Eq. (9)) is discretized as
@@
and $\lambda$ is the interfacial coefficient appearing in the weighted jump definition Eq. (6) (e.g. $\lambda\equiv 1$ for continuity).
@@ 4.3 Discrete equations
We now present a $\theta$-scheme time discretization of the semi-discrete bulk balances (25). For $0\leq\theta\leq 1$, we approximate fluxes and sources at $t_{n+\theta}:=(1-\theta)t_{n}+\theta t_{n+1}$:
@@
and $\Delta t_{n}:=t_{n+1}-t_{n}$. The choices $\theta=0$ and $\theta=1$ correspond to forward and backward Euler, respectively and $\theta=\tfrac{1}{2}$ to the midpoint rule.
@@ 4.4 Block structure of the linear system
Since the continuum model is linear, the corresponding fully discrete equations are linear and can be written in matrix-vector form. Even for $\theta=0$ (forward Euler), the unknowns at time level $n+1$ remain coupled through the interfacial conditions. The matrices are nevertheless sparse, as the discrete operators were designed to keep the stencil compact. The numerical solution at each step therefore amounts to solving a sparse linear system coupling bulk and interfacial unknowns.
We first eliminate the face fluxes $Q^{\alpha\pm}$ using the discrete constitutive relation (39), leaving $\Phi^{\omega\pm}$ and $\Phi^{\gamma\pm}$ as the remaining unknowns. Let the (diagonal) mass matrices be defined by
@@
Let $\operatorname{div}^{\alpha}$ denote the full (bulk+interface) directional divergence operator used in (38) and let $\operatorname{grad}^{\alpha\omega}$ and $\operatorname{grad}^{\alpha\gamma}$ denote the bulk and interfacial contributions to the directional gradient. Since these operators are linear in their arguments, we introduce the linear operators
$L^{\omega\omega,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\bigl(V^{\pm}\,\operatorname{div}^{\alpha}(B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\operatorname{grad}^{\alpha\omega}(B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (42)
$L^{\omega\gamma,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\bigl(V^{\pm}\,\operatorname{div}^{\alpha}(B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\operatorname{grad}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (43)
$L^{\gamma\omega,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\bigl(V^{\pm}\,\operatorname{div}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\operatorname{grad}^{\alpha\omega}(B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (44)
$L^{\gamma\gamma,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\bigl(V^{\pm}\,\operatorname{div}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\operatorname{grad}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},W^{\alpha\pm},\,\cdot).$ (45)
@@
The fully discrete system at a given time step can then be written as
@@
Finally, for symmetric positive definite face-centered diffusion coefficients $K^{\alpha\pm}$, the operators $L^{\omega\omega,\pm}$ are symmetric positive (semi-)definite on the active bulk unknowns. Moreover, under the standard finite-volume duality between divergence and gradient, one typically has the discrete adjoint relation
@@
The resulting sparse linear system (46) is solved using an appropriate linear solver (direct or iterative), depending on problem size. A one-dimensional example is presented in App. C.
@@ 5 Numerical validation in static domains
We validate the accuracy and robustness of the proposed cut-cell formulation for static embedded geometries. Both monophasic and diphasic configurations are considered, for elliptic (steady diffusion) and parabolic (unsteady diffusion) problems. Analytical solutions and corresponding forcing terms used for verification are provided in each paragraph.
@@
5.1 Monophasic validation
We first assess the monophasic formulation in a static embedded geometry. The monophasic formulation is obtained by restricting the computation to the light phase $\Omega^{-}$, i.e. by solving only for $\Phi^{\omega-}$ and $\Phi^{\gamma-}$ and dropping all “++” variables and interfacial coupling equations. In the continuous setting, the embedded boundary $\Gamma=\partial\Omega^{-}\cap\Omega$ is equipped with the Robin condition
@@ 5.1 Monophasic validation
$\sum_{\alpha=1}^{d}V^{-}_{i,j}\,\operatorname{div}^{\alpha\gamma}_{i,j}\!\left(A^{\alpha-},B^{\alpha-},Q^{\alpha-}_{n+\theta}\right)\;+\;\beta\,\langle\Gamma_{i,j},1\rangle\,\Phi^{\gamma-}_{n+\theta,i,j}\;=\;\langle\Gamma_{i,j},1\rangle\,G_{n+\theta,i,j},$
@@ 5.1.1 Johansen-Colella Problem 1: star-shaped Poisson problem
The right-hand side in (50) follows from the polar Laplacian identity $\Delta(r^{m}\cos(k\theta))=(m^{2}-k^{2})\,r^{m-2}\cos(k\theta)$ with $m=4$ and $k=3$.
The problem is discretized on a uniform Cartesian grid of spacing $h$ using the cut-cell operators of Sec. 4, with $\Gamma$ embedded in the background mesh. Errors are reported using the $L^{2}$ norms defined in Sec. 5, distinguishing regular cells (fully inside $\Omega$), cut cells (intersected by $\Gamma$) and all active cells.
@@
Figure 10: Johansen-Colella Problem 1 on the star-shaped domain (49): numerical solution and corresponding $\log_{10}$ error field on a $512^{2}$ mesh
@@
By the maximum principle, the exact solution is bounded within $[0,1]$. We therefore monitor the discrete extrema $\min\phi^{h}$ and $\max\phi^{h}$ and the overshoot/undershoot ratio $k/N$, where $k$ is the number of active cells with values outside $[0,1]$ and $N$ is the total number of active cells. No overshoot or undershoot is observed at any resolution, as summarized in Table 2.
Figure 11: Johansen-Colella Problem 2: numerical steady diffusion field $\phi^{\rm num}$ in the unit box, with the embedded flower-shaped boundary (52) defining a hole.
@@
We next validate the monophasic formulation for an elliptic problem with a Robin boundary condition imposed on a curved embedded boundary. The computational box is $\Omega_{\rm box}=[0,4]\times[0,4],$ and the physical domain is the disk
@@
which satisfies $\Delta\phi^{\rm ex}=1$ in $\Omega$ and the Robin condition (56) on $\Gamma$.
@@
Figure 12: Poisson problem in a disk with Robin boundary condition (56): numerical solution and base-10 logarithm of the pointwise error.
Table 3 reports the $L^{2}$ errors evaluated over all, regular and cut cells, together with the $H^{1}$ error restricted to cut cells. The discrete $H^{1}$ error is defined from the gradient of the numerical solution,
@@
$\|e\|_{H^{1},\mathrm{cut}}=\left(\frac{\displaystyle\sum_{(i,j)\in\mathcal{I}_{\mathrm{cut}}}V_{i,j}\,\bigl\|\nabla\Phi_{i,j}-\nabla\Phi^{\rm ex}_{i,j}\bigr\|^{2}}{\displaystyle\sum_{(i,j)\in\mathcal{I}_{\mathrm{cut}}}V_{i,j}\,}\right)^{1/2},$
@@
where $\nabla\Phi$ is obtained with the discrete gradient operator of Sec. 4.1 and $\nabla\Phi^{\rm ex}$ is evaluated at consistent locations. Pairwise orders are computed using (48).
$h$ $N_{x}{=}N_{y}$ $\|e\|_{2,\mathrm{all}}$ $\|e\|_{2,\mathrm{reg}}$ $\|e\|_{2,\mathrm{cut}}$ $p_{2,\mathrm{all}}$ $p_{2,\mathrm{reg}}$ $p_{2,\mathrm{cut}}$ $\|e\|_{H^{1},\mathrm{cut}}$ $p_{H^{1},\mathrm{cut}}$
0.25 8 5.413e35.413e{-3} 3.883e33.883e{-3} 3.771e33.771e{-3} - - - 4.578e24.578e{-2} -
0.125 16 1.216e31.216e{-3} 8.813e48.813e{-4} 8.382e48.382e{-4} 2.15 2.14 2.17 1.738e21.738e{-2} 1.40
0.0625 32 4.198e44.198e{-4} 2.616e42.616e{-4} 3.284e43.284e{-4} 1.53 1.75 1.35 8.419e38.419e{-3} 1.05
0.03125 64 1.209e41.209e{-4} 7.104e57.104e{-5} 9.788e59.788e{-5} 1.80 1.88 1.75 4.550e34.550e{-3} 0.89
0.015625 128 3.476e53.476e{-5} 1.697e51.697e{-5} 3.033e53.033e{-5} 1.80 2.07 1.69 1.871e31.871e{-3} 1.28
0.0078125 256 1.000e51.000e{-5} 4.056e64.056e{-6} 9.142e69.142e{-6} 1.80 2.06 1.73 9.086e49.086e{-4} 1.04
@@
Table 3: Poisson problem in a disk with Robin boundary condition: $L^{2}$ errors over all, regular and cut cells and $H^{1}$ error restricted to cut cells, with pairwise convergence orders.
@@
Table 3 confirms that the cut-cell formulation enforces the Robin condition on the curved embedded boundary with the expected accuracy. The $L^{2}$ error over regular cells converges at essentially second order once the mesh is sufficiently fine ($p_{2,\mathrm{reg}}\approx 2$ for the last refinements), indicating that the scheme retains its nominal accuracy away from the boundary. Over all cells, the observed rate stabilizes around $p_{2,\mathrm{all}}\simeq 1.8$, reflecting the fact that the global error is influenced by the cut-cell region where geometric trimming and boundary closure dominate. The $L^{2}$ error restricted to cut cells converges slightly below second order ($p_{2,\mathrm{cut}}\approx 1.7$ on the finest grids), which is consistent with the reduced regularity of the discrete operators near embedded boundaries. Finally, the cut-cell $H^{1}$ error exhibits approximately first-order convergence ($p_{H^{1},\mathrm{cut}}\approx 1$), as expected for gradient reconstruction in the vicinity of irregular control volumes and curved boundary closures.
A key feature of the proposed cut-cell formulation is that boundary and interface conditions are enforced directly at the discrete level. As a result, interfacial unknowns $\Phi_{i,j}^{\gamma}$ are part of the primary solution vector and provide direct access to the boundary/interface values without any post-processing or interpolation from nearby cell averages. This is illustrated in Fig. 13, which reports the interfacial values of $\phi$ along the polar coordinate and shows that the relative error at the interface remains below 0.33%
Figure 13: Poisson problem with Robin boundary condition: interfacial values of $\phi$ as a function of the polar angle along the circular embedded boundary. The analytical solution is shown in red.
@@
We finally consider a three-dimensional monophasic parabolic test in a spherical domain with a Robin boundary condition. The physical domain is the ball $\Omega=\{\mathbf{x}\in\mathbb{R}^{3}:\ \|\mathbf{x}\|<R\},$ embedded in a fixed Cartesian box. Let $r=\|\mathbf{x}\|$. We solve
@@
$\phi(r,t)=\sum_{n=1}^{\infty}C_{n}\,\frac{\sin(\mu_{n}r/R)}{r}\,\exp\!\left(-a\mu_{n}^{2}\,\frac{t}{R^{2}}\right),$ (61)
@@
Table 4: $L^{2}$ errors at final time $t_{f}=0.1$ for the 3D monophasic unsteady diffusion problem in a sphere with Robin boundary condition.
Table 4 demonstrates that the proposed cut-cell framework extends cleanly to three dimensions: the geometric moments (trimmed volumes, face apertures and boundary measures) can be integrated robustly in 3D and lead to stable discretizations even on very coarse grids (e.g. $N_{\rm diam}=1$-3). On the finest refinements, the global error converges at approximately second order, with fitted rates $p_{\mathrm{all}}\approx 1.9$ and $p_{\mathrm{reg}}\approx 1.8$. The cut-cell error also exhibits near-second-order behavior ($p_{\mathrm{cut}}\approx 2.0$ in the fit), indicating that the Robin closure and the moment-based representation of the curved boundary remain accurate in three dimensions.
@@
Fig. 14 shows the radial profile of $\phi(\cdot,t_{f})$ extracted along a line passing through the center. Blue markers denote the numerical solution sampled at cell centroids as a function of the radius $r=\|\mathbf{x}\|$, while the red marker highlights the discrete boundary value at $r=R$ (the cut-cell boundary/interfacial unknown used to impose (59)). The close agreement with the analytical radial series (61) confirms that the moment-weighted boundary closure remains accurate in 3D and that the boundary degree of freedom provides a consistent trace of the solution at $\Gamma$.
Figure 14: 3D unsteady diffusion in a sphere with Robin boundary condition: radial profile of $\phi(r,t_{f})$ along a line through the center. Blue markers: numerical samples as a function of $r$. Red marker: discrete boundary/interfacial value at $r=R$ used in the Robin closure. Solid line: analytical series solution (61).
Neumann limit and mass conservation.
To verify mass conservation in the pure Neumann limit, we set, $\partial_{n}\phi=0\qquad\text{on }\Gamma.$ The analytical configuration is chosen such that the total mass is constant in time. On the discrete level, we monitor the numerical mass and the relative drift,
$M^{h}(t)=\sum_{\text{active cells }i}V_{i}\,\phi_{i}(t),\quad\delta M^{h}(t)=\frac{M^{h}(t)-M^{h}(0)}{M^{h}(0)}.$
Using the same time-integration parameters as in the Robin tests (midpoint rule, $t_{f}=0.1$), the recorded mass remains constant up to machine precision. A representative run gives the values reported in Table 5.
$t$ $M^{h}(t)$ $\delta M^{h}(t)$ $|\delta M^{h}(t)|_{\max,\,0\leq t\leq t_{f}}$
0 $1.9635\times 10^{-1}$ 0 $3.1\times 10^{-17}$
$t_{f}/2$ $1.9635\times 10^{-1}$ 0
$t_{f}$ $1.9635\times 10^{-1}$ $2.8\times 10^{-17}$
Table 5: Neumann test: discrete mass $M^{h}(t)$ and relative drift $\delta M^{h}(t)$ over time. The total mass is preserved to within $\mathcal{O}(10^{-17})$, confirming exact conservation of the discrete Neumann fluxes in the cut-cell formulation.
5.2 Diphasic validation

### v2

@@ 4.1.2 Gradient operator
The interfacial contribution (last term in Eq. (32)) is approximated by
@@
Finally, the constitutive relation Eq. (4) is imposed in discrete form by
@@ 4.2 Semi-discrete interface conditions
The interfacial conservation law Eq. (7) corresponds to continuity of the diffusive flux,
@@
The integrals in Eq. (35) correspond to the interfacial contributions appearing in the Gauss decomposition in Eq. (26). They are approximated using the discrete interfacial divergence operators $\operatorname{div}^{\alpha\gamma}$ introduced in Eq. (30). Consequently, the flux balance is semi-discretized as
@@
For mixed cells $(i,j)$, the jump/closure condition in Eq. (8) (or in Eq. (9)) is semi-discretized as
@@
and $\lambda$ is the interfacial coefficient appearing in the weighted jump definition Eq. (6) (e.g. $\lambda=1$ adn $f=0$ to impose continuity of $\phi$ across the interface).
@@ 4.3 Discrete equations
We now use a $\theta$-scheme time discretization of the semi-discrete bulk balances Eq. (25). For $0\leq\theta\leq 1$, we approximate fluxes and sources at $t_{n+\theta}:=(1-\theta)t_{n}+\theta t_{n+1}$:
@@
and $\Delta t_{n}:=t_{n+1}-t_{n}$. The choices $\theta=0$ and $\theta=1$ correspond to the forward and backward Euler schemes respectively and $\theta=\tfrac{1}{2}$ to the midpoint rule.
@@ 4.4 Block structure of the linear system
Since the continuum model is linear, the corresponding discrete equations are linear and can be written in matrix-vector form. Even for $\theta=0$ (forward Euler), the unknowns at time level $n+1$ remain coupled through the interfacial conditions. The matrices are nevertheless sparse, as the discrete operators were designed to keep the stencil compact. The numerical solution at each step therefore amounts to solving a sparse linear system coupling bulk and interfacial unknowns.
We first eliminate the face fluxes $Q^{\alpha\pm}$ using the discrete constitutive relation Eq. (39), leaving $\Phi^{\omega\pm}$ and $\Phi^{\gamma\pm}$ as the remaining unknowns. Let the (diagonal) mass matrices be defined by
@@
Let $\operatorname{div}^{\alpha}$ denote the full (bulk+interface) directional divergence operator used in Eq. (38) and let $\operatorname{grad}^{\alpha\omega}$ and $\operatorname{grad}^{\alpha\gamma}$ denote the bulk and interfacial contributions to the directional gradient. Since these operators are linear in their arguments, we introduce the linear operators
$L^{\omega\omega,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\partial_{Q}\bigl(V^{\pm}\,\operatorname{div}^{\alpha}(B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\partial_{\Phi}\operatorname{grad}^{\alpha\omega}(B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (42)
$L^{\omega\gamma,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\partial_{Q}\bigl(V^{\pm}\,\operatorname{div}^{\alpha}(B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\partial_{\Phi}\operatorname{grad}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (43)
$L^{\gamma\omega,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\partial_{Q}\bigl(V^{\pm}\,\operatorname{div}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\partial_{\Phi}\operatorname{grad}^{\alpha\omega}(B^{\alpha\pm},W^{\alpha\pm},\,\cdot),$ (44)
$L^{\gamma\gamma,\pm}$ $:=-\sum_{\alpha\in\{1,2\}}\partial_{Q}\bigl(V^{\pm}\,\operatorname{div}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},\,\cdot)\bigr)\;K^{\alpha\pm}\;\partial_{\Phi}\operatorname{grad}^{\alpha\gamma}(A^{\alpha\pm},B^{\alpha\pm},W^{\alpha\pm},\,\cdot).$ (45)
@@
The discrete system at a given time step can then be written as
@@
Finally, for symmetric positive definite face-centered diffusion coefficients $K^{\alpha\pm}$, the operators $L^{\omega\omega,\pm}$ are symmetric positive definite on the active bulk unknowns. Moreover, under the standard finite-volume duality between divergence and gradient, the discrete adjoint relation also holds,
@@
The resulting sparse linear system Eq. (46) is solved using an appropriate linear solver (direct or iterative), depending on problem size. A one-dimensional example is presented in App. C.
@@ 5 Numerical validation in static domains
We validate the accuracy and robustness of the proposed cut-cell formulation for static embedded geometries. Both single-phase and two-phase configurations are considered, for elliptic (steady diffusion) and parabolic (unsteady diffusion) problems. Analytical solutions and corresponding forcing terms used for verification are provided in each paragraph.
@@
In addition, a global convergence rate is computed as the least-squares slope of $\log(\|e\|_{2,S})$ versus $\log(h)$ over all considered grid resolutions.
5.1 Single-phase validation
We first assess the single-phase formulation in a static embedded geometry. The single-phase formulation is obtained by restricting the computation to the light phase $\Omega^{-}$, i.e. by solving only for $\Phi^{\omega-}$ and $\Phi^{\gamma-}$ and dropping all “++” variables and interfacial coupling equations. In the continuous setting, the embedded boundary $\Gamma=\partial\Omega^{-}\cap\Omega$ is equipped with the Robin condition
@@ 5.1 Single-phase validation
$\sum_{\alpha=1}^{d}V^{-}_{i,j}\,\operatorname{div}^{\alpha\gamma}_{i,j}\!\left(A^{\alpha-},B^{\alpha-},Q^{\alpha-}_{n+\theta}\right)\;+\;\beta\,\langle\Gamma_{i,j},1\rangle\,\Phi^{\gamma-}_{n+\theta,i,j}\;=\;\langle\Gamma_{i,j},1\rangle\,F_{n+\theta,i,j},$
@@ 5.1.1 Johansen-Colella Problem 1: Star-shaped Poisson problem
The right-hand side in Eq. (50) follows from the polar Laplacian identity $\Delta(r^{m}\cos(k\theta))=(m^{2}-k^{2})\,r^{m-2}\cos(k\theta)$ with $m=4$ and $k=3$.
The problem is discretized on a uniform Cartesian grid of spacing $h$ using the cut-cell operators of Sec. 4, with $\Gamma$ embedded in the background mesh. Errors are reported using the $L^{2}$ norms defined in Eq. (47), distinguishing regular cells (fully inside $\Omega$), cut cells (intersected by $\Gamma$) and all active cells.
@@
Figure 10: Johansen-Colella Problem 1 on the star-shaped domain Eq. (49): numerical solution and corresponding $\log_{10}$ error field on a $512^{2}$ mesh
@@
By the maximum principle, the exact solution is bounded within $[0,1]$. We therefore monitor the discrete extrema $\min\phi^{h}$ and $\max\phi^{h}$ and the overshoot/undershoot ratio $k/N$, where $k$ is the number of active cells with values outside $[0,1]$ and $N$ is the total number of active cells. No overshoot or undershoot is observed at any resolution, as summarized in Tab. 2. To stress the robustness of the proposed method, extremely coarse resolution are also considered here (only $4\times 4$ cells).
Figure 11: Johansen-Colella Problem 2: numerical steady diffusion field $\phi^{\rm num}$ in the unit box, with the embedded flower-shaped boundary Eq. (52) defining a hole.
@@
We next validate the single-phase formulation for an elliptic problem with a Robin boundary condition imposed on a curved embedded boundary. The computational box is $\Omega_{\rm box}=[0,4]\times[0,4],$ and the physical domain is the disk
@@
which satisfies $\Delta\phi^{\rm ex}=1$ in $\Omega$ and the Robin condition Eq. (56) on $\Gamma$.
@@
Figure 12: Poisson problem in a disk with Robin boundary condition Eq. (56): numerical solution and base-10 logarithm of the pointwise error.
The discrete semi-norm $H^{1}$ error is defined from the gradient of the numerical solution,
@@
$\|e\|_{H^{1},{S}}=\left(\frac{\displaystyle\sum_{(i,j)\in\mathcal{I}_{S}}V_{i,j}\,\bigl\|\nabla\Phi_{i,j}-\nabla\Phi^{\rm ex}_{i,j}\bigr\|^{2}}{\displaystyle\sum_{(i,j)\in\mathcal{I}_{S}}V_{i,j}\,}\right)^{1/2},$
@@
where $\nabla\Phi$ is obtained with the discrete gradient operator of Sec. 4.1 and $\nabla\Phi^{\rm ex}$ is evaluated at consistent locations. Pairwise orders are computed using Eq. (48). Tab. 3 reports the $L^{2}$ errors evaluated over all, regular and cut cells. Tab. 4 reports the $H^{1}$ semi-norm errors evaluated over all, regular and cut cells.
$h$ $N_{x}{=}N_{y}$ $\|e\|_{2,\mathrm{all}}$ $\|e\|_{2,\mathrm{reg}}$ $\|e\|_{2,\mathrm{cut}}$ $p_{2,\mathrm{all}}$ $p_{2,\mathrm{reg}}$ $p_{2,\mathrm{cut}}$
0.25 8 5.413e35.413e{-3} 3.883e33.883e{-3} 3.771e33.771e{-3} - - -
0.125 16 1.216e31.216e{-3} 8.813e48.813e{-4} 8.382e48.382e{-4} 2.15 2.14 2.17
0.0625 32 4.198e44.198e{-4} 2.616e42.616e{-4} 3.284e43.284e{-4} 1.53 1.75 1.35
0.03125 64 1.209e41.209e{-4} 7.104e57.104e{-5} 9.788e59.788e{-5} 1.80 1.88 1.75
0.015625 128 3.476e53.476e{-5} 1.697e51.697e{-5} 3.033e53.033e{-5} 1.80 2.07 1.69
0.0078125 256 1.000e51.000e{-5} 4.056e64.056e{-6} 9.142e69.142e{-6} 1.80 2.06 1.73
fit - - - - 1.79 1.96 1.70
@@
Table 3: $L^{2}$ error convergence for the Poisson problem in a disk with Robin boundary condition, measured over all, regular, and cut cells.
$h$ $N_{x}{=}N_{y}$ $\|e\|_{H^{1},\mathrm{all}}$ $\|e\|_{H^{1},\mathrm{reg}}$ $\|e\|_{H^{1},\mathrm{cut}}$ $p_{H^{1},\mathrm{all}}$ $p_{H^{1},\mathrm{reg}}$ $p_{H^{1},\mathrm{cut}}$
0.25 8 8.888e28.888e{-2} 7.618e27.618e{-2} 4.578e24.578e{-2} - - -
0.125 16 4.456e24.456e{-2} 4.103e24.103e{-2} 1.738e21.738e{-2} 1.00 0.89 1.40
0.0625 32 2.290e22.290e{-2} 2.130e22.130e{-2} 8.419e38.419e{-3} 0.96 0.95 1.05
0.03125 64 1.181e21.181e{-2} 1.089e21.089e{-2} 4.550e34.550e{-3} 0.96 0.97 0.89
0.015625 128 5.797e35.797e{-3} 5.487e35.487e{-3} 1.871e31.871e{-3} 1.03 0.99 1.28
0.0078125 256 2.909e32.909e{-3} 2.763e32.763e{-3} 9.086e49.086e{-4} 0.99 0.99 1.04
fit - - - - 0.98 0.96 1.11
Table 4: $H^{1}$-semi-norm error convergence for the Poisson problem in a disk with Robin boundary condition, measured over all, regular, and cut cells.
@@
Tab. 3 confirms that the cut-cell formulation enforces the Robin condition on the curved embedded boundary with the expected accuracy. The $L^{2}$ error over regular cells converges at essentially second order once the mesh is sufficiently fine ($p_{2,\mathrm{reg}}\approx 2$ for the last refinements), indicating that the scheme retains its nominal accuracy away from the boundary. Over all cells, the observed rate stabilizes around $p_{2,\mathrm{all}}\approx 1.8$, reflecting the fact that the global error is influenced by the cut-cell region where geometric trimming and boundary closure dominate. The $L^{2}$ error restricted to cut cells converges slightly below second order ($p_{2,\mathrm{cut}}\approx 1.7$ on the finest grids), which is consistent with the reduced regularity of the discrete operators near embedded boundaries. Finally, Tab. 4 shows that the discrete $H^{1}$-semi-norm errors converge at first order, with overall slopes close to one ($p_{H^{1},\mathrm{reg}}\approx 1$, $p_{H^{1},\mathrm{all}}\approx 1$ and $p_{H^{1},\mathrm{cut}}\approx 1$), as expected for gradient reconstruction in the vicinity of irregular control volumes and curved boundary closures.
@@
We finally consider a three-dimensional single-phase parabolic test in a spherical domain with a Robin boundary condition. The physical domain is the ball $\Omega=\{\mathbf{x}\in\mathbb{R}^{3}:\ \|\mathbf{x}\|<R\},$ embedded in a fixed Cartesian box. Let $r=\|\mathbf{x}\|$. We solve
@@
$\phi(r,t)=\sum_{n=1}^{\infty}C_{n}\,\frac{\sin(\mu_{n}r/R)}{r}\,\exp\!\left(-a\mu_{n}^{2}t/R^{2}\right),$ (61)
@@
Table 5: $L^{2}$ errors at final time $t_{f}=0.1$ for the 3D single-phase unsteady diffusion problem in a sphere with Robin boundary condition.
Tab. 5 demonstrates that the proposed cut-cell framework extends to three dimensions: the geometric moments (trimmed volumes, face apertures and boundary measures) can be integrated robustly in 3D and lead to stable discretizations even on very coarse grids (e.g. $N_{\rm diam}=1$). On the finest refinements, the global error converges at approximately second order, with fitted rates $p_{\mathrm{all}}\approx 1.9$ and $p_{\mathrm{reg}}\approx 1.8$. The cut-cell error also exhibits near-second-order behavior ($p_{\mathrm{cut}}\approx 2.0$ ), indicating that the Robin closure and the moment-based representation of the curved boundary remain accurate in three dimensions.
@@
Fig. 13 shows the radial profile of $\phi(\cdot,t_{f})$ extracted along a line passing through the center. Blue markers denote the numerical solution sampled at cell centroids as a function of the radius $r=\|\mathbf{x}\|$, while the red marker highlights the discrete boundary value at $r=R$ (the cut-cell boundary/interfacial unknown used to impose Eq. (59)). The close agreement with the analytical radial series Eq. (61) confirms that the moment-weighted boundary closure remains accurate in 3D and that the boundary degree of freedom provides a consistent trace of the solution at $\Gamma$.
Figure 13: 3D unsteady diffusion in a sphere with Robin boundary condition: radial profile of $\phi(r,t_{f})$ along a line through the center. Blue markers: numerical samples as a function of $r$. Red marker: discrete boundary/interfacial value at $r=R$ used in the Robin closure. Solid line: analytical series solution Eq. (61).
Neumann limit and global conservation.
To verify global conservation in the pure Neumann limit, we set $\partial_{n}\phi=0\text{ on }\Gamma.$ The analytical configuration is chosen such that the domain-integrated quantity is constant in time. On the discrete level, we monitor the conserved integral and its relative drift,
$I^{h}(t)=\sum_{\text{active cells }i}V_{i}\,\phi_{i}(t),\quad\delta I^{h}(t)=\frac{I^{h}(t)-I^{h}(0)}{I^{h}(0)}.$
Using the same time-integration parameters as in the Robin tests (midpoint rule, $t_{f}=0.1$), $I^{h}(t)$ remains constant up to machine precision: the maximum drift over $0,t_{f}]$ is $|\delta I^{h}|_{\max}=3.1\times 10^{-17}$, and at the final time we obtain $\delta I^{h}(t_{f})=2.8\times 10^{-17}$.
5.2 Two-phase validation